OpenStudy (anonymous):

-2x+5y=8 find the value of y when x equals -14

OpenStudy (anonymous):

-2x + 5y = 8 -2(-14) + 5y = 8 28 + 5y = 8 5y = 8 - 28 5y = -20 y = -4

OpenStudy (anonymous):

thx again